### v3

# Shallow learning enables real-time inference of molecular composition from spectroscopy of brain tissue

Abstract
Optical imaging modalities such as near-infrared spectroscopy (NIRS) and hyperspectral imaging (HSI) represent a promising alternative for low-cost, non-invasive, and fast monitoring of functional and structural properties of living tissue. Particularly, the possibility of extracting the molecular composition of the tissue from the optical spectra in real-time deems the spectroscopy techniques as a unique diagnostic tool. However, due to a lack of paired optical and molecular profiling studies, building a mapping between a spectral signature and a corresponding set of molecular concentrations is still an unsolved problem. Furthermore, there are currently no established methods to streamline inference of the biochemical composition from the optical spectrum for real-time applications such as surgical monitoring.
In this paper, we develop a technique for fast inference of changes in the molecular composition of brain tissue. We base our method on the Beer-Lambert law to analytically connect the spectra with concentrations and use a deep-learning approach to significantly speed up the concentration inference compared to traditional optimization methods. We test our approach on real data obtained from the broadband NIRS study of piglets’ brains and the HSI dataset of glioma patients. The results demonstrate that the proposed method enables real-time molecular composition inference while maintaining the accuracy of traditional linear and non-linear optimization solvers.
keywords:
optics, photonics, light, hyperspectral, machine learning, neural networks, brain imaging
[…]
normal-$\dagger$ The authors contributed equally
1 Introduction
Various biomedical applications such as laboratory tissue examination or neurosurgery require access to rapid monitoring of intrinsic tissue properties. In particular, neuronavigation would benefit by having structural and functional information about brain matter in real-time [1, 2, 3]. Spatially-resolved maps of the tissue characteristics would allow bypassing invasive disease diagnostics, e.g., biopsy, which halt the operation. Instead, a surgical decision could be made during the operation, reducing its time and preserving a healthy brain.
Optical sensing, such as near-infrared spectroscopy (NIRS) and hyperspectral imaging (HSI), emerge as promising alternatives to address these clinical needs [4, 5]. These techniques can probe biological matter utilizing non-ionizing electromagnetic radiation within the visible and near-infrared ranges. The spectral instrumentation is inexpensive, portable, and allows for continuous monitoring at the bedside [6].
The overarching principle behind spectroscopy-based molecular characterization is to relate the reflection spectrum obtained upon illumination of the tissue surface with its optical properties. The molecules constituting the tissue have unique absorption dependency on the radiation frequency, and thus, the reflection should exhibit molecular absorption signatures in its frequency dependency.
However, several other physical factors contribute to shaping the spectra. These include light scattering on the surface and within tissue volume, autofluorescence, and background illumination [7]. Disentangling these phenomena from a reflectance spectrum is often an ambiguous, ill-posed problem, yet it is crucial for deducing the relation between the reflection and molecular composition. Another complication is a scarcity of available studies in which optical monitoring and biochemical composition analysis are performed simultaneously.
Analytical and statistical approaches exist aiming to unmix optical spectra into the physical phenomena defining the spectra profile under a limited data regime [8, 9, 10, 11, 12, 13]. The first family of methods mitigates the data scarcity by introducing a physical prior to establish the spectrochemical link. Typically, the Beer-Lambert law [14, 5] is used to provide such a link by describing the incoming light’s energy dissipation as an exponentially decaying function of light penetration depth into the tissue:
$\log(I_{R}(\lambda)/I_{0}(\lambda))=-\left[\sum_{i}{c_{i}\mu^{i}_{a}(\lambda)}+s\mu_{s}(\lambda)\right]\cdot l+U$ (1)
Here, $I_{R}(\lambda)$ and $I_{0}(\lambda)$ are the intensities of the reflected and the incoming light, $\mu_{a}$ and $\mu_{s}$ is the absorption and scattering coefficients, the index $i$ denotes the molecule constituting the tissue such as water, fat, hemoglobin or cytochromes, $c_{i}$ denotes the corresponding concentration, and $s$ is the weight of scattering in the total light energy dissipation. The remaining quantities are $\lambda$, which is the light wavelength, $l$ is the light penetration depth, and $U$ describes other physical factors contributing to the energy dissipation of the incoming light or other sources of the optical signal captured by the cameras. Now, in the case of changes in molecular composition over the course of optical monitoring, one can assume that the effects contributing to $U$ either stay constant (e.g., which is a fair assumption for ambient illumination) or change notably less than the total absorption  [^1] In certain scenarios, even changes in scattering are not expected since it is a rather bulk effect dependent on the density of the probed matter rather than molecule-specific one.. Under this assumption, subtraction of two reflection spectra, $\log I_{R}^{2}-\log I_{R}^{1}$, would cancel out or make negligible the term $\delta U=U^{2}-U^{1}$ in the following equation:
$\log(I_{R}^{2}(\lambda)/I_{R}^{1}(\lambda))=-\left[\sum_{i}{\delta c_{i}\mu^{i}_{a}(\lambda)}+\delta(s\mu_{s}(\lambda))\right]\cdot l+\delta U$ (2)
In such a differential form, the Beer-Lambert law can now be used to identify molecular composition. For this, standard least-squares optimization algorithms can be employed to minimize the difference between the real spectra and the spectra obtained from the Beer-Lambert law. As a result of the minimization, the optimal values of the set of concentration changes $\{\delta c_{i}=c_{i}^{2}-c_{i}^{1}\}$ are obtained (alongside the scattering parameters).
Another family of unmixing approaches is based on statistical techniques such as principal component analysis (PCA) [12, 13, 11]. Here, by analyzing a large dataset of reflection spectra using the PCA, one identifies the importance of each wavelength to explain the variance in the dataset. Assuming that from the physics point of view, the variance in the reflection spectra is mainly due to varying molecular absorption, one can try to correlate the output of the PCA analysis with the absorption spectra. In [11], the authors discuss how such correlation can reveal the relative difference in concentration of chromophores.
The overarching drawback of these approaches is the computational time it takes to infer the biochemical composition. For example, the optimization methods take a subsecond time to infer the composition of a single spectrum containing typical for NIRS and HSI number of wavelengths (a few hundred). However, when using modalities providing high spatial resolution imaging for real-time applications, one needs to solve the optimization task in a subsecond time for as many spectra as there are spatial pixels, as every pixel contains its own spectrum. For modalities like HSI, the number of pixels can be easily in the order of $10^{5}$. Providing subsecond timings for simultaneous inference on such an amount of spectra poses a challenge for traditional methods  [^2] Note that under the assumption of negligible contribution from scattering in the modified Beer-Lambert law, the Eq. 2 results in a linear system. Such a system can be efficiently solved using the pseudoinverse, achieving close to real-time computation. However, in general, when including non-linear terms in the Beer-Lambert law formulation (scattering, non-linear absorption effects, etc.), the system of equations can not be solved anymore using pseudoinverse..
The present paper aims to alleviate the computational time limitation of existing methods by proposing a data-driven solution that significantly speeds up the biochemical composition inference. We test two types of deep-learning-based approaches which differ in the way the training dataset is collected: a) directly by generating synthetic spectrum-concentrations pairs from the Beer-lambert law or b) by inferring the molecular composition from real spectra by means of the traditional optimization and then use the pairs of real spectra and inferred concentrations for a neural network training. We evaluate the method on broadband NIRS data (transmission mode) [15] and hyperspectral data (reflection mode) [16] of brain tissue. We achieve a significant speed-up compared to the convex least-squares optimization while maintaining accuracy. To our knowledge, this is the first work that applies a neural-network-based approach to provide real-time inference of cerebral chromophore composition from spectroscopy measurements [17, 18, 19, 20, 21, 22, 23, 24, 25].
Figure 1: The general pipeline describing our approach for inferring molecular concentrations’ changes $\{\delta c_{i}\}$ from changes in the reflection spectra.
2 Method
As mentioned in the introduction, inference of absolute chromophore concentrations from an optical spectrum is a challenging task due to multiple physical effects shaping the reflection spectrum. Thus, we instead aim to predict the changes in concentrations from changes in the spectra, Eq. 2.
Datasets creation.
In our method, we use a supervised data-driven approach by creating a dataset of attenuation-concentration pairs to train a neural network (by attenuation, we imply the logarithm of the reflection: $\Delta A=\log I_{R}^{2}-\log I_{R}^{1}$). We employ two different strategies to create the dataset:
a) The first strategy directly utilizes the Beer-Lambert law to generate the training dataset pairs $(\Delta A,\{\delta c_{i}\})$ with $\Delta A$ being the difference in attenuation between two spectra and $\{\delta c_{i}\}$ the corresponding differences in concentration of chromophores. For each chromophore, we randomly sample values for changes in the molecular composition $\{\delta c_{i}\}$ using the uniform distribution within physiologically plausible ranges.
If scattering is included in Eq. 2, we may assume it to be of rational form:
$s\mu_{s}(\lambda)=s\left(\frac{\lambda}{500nm}\right)^{-b}$ (3)
with the scaling of the anisotropy $g=0.9$ included in $s=s^{\prime}/(1-g)$ [26, 5]. In the differential form, we obtain
$\delta(s\mu_{s}(\lambda))=s_{2}\left(\frac{\lambda}{500nm}\right)^{-b_{2}}-s_{1}\left(\frac{\lambda}{500nm}\right)^{-b_{1}}$ (4)
and therefore also uniformly sample parameters $s_{1},s_{2},b_{1},b_{2}$ within plausible ranges [26]. Subsequently, we input the obtained values into the Beer-Lambert law to obtain synthetic differential attenuations $\Delta A(\lambda)$. This difference in attenuation $\Delta A$ as input and the corresponding $\{\delta c_{i}\}$ as output are then used for training.
Figure 2: Two strategies for collecting the training dataset. Strategy (a) in which we train a network on synthetic attenuation-concentration pairs generated from the modified Beer-Lambert law. Strategy (b) in which the training is performed on pairs of real attenuation and concentrations obtained through the least-squares fit to the corresponding real spectra.
b) Given that the distribution of the synthetic spectra obtained according to the strategy described above can be notably different from the distribution of real spectra, this can result in an unsatisfactory network prediction accuracy. Therefore, in addition to (a), we evaluate another strategy for creating a dataset trying to bridge the gap between the physical model and real data. For this, we use traditional least squares minimization to fit the changes in the real reflectance spectrum with the modified Beer-Lambert law. The concentrations $\{\delta c_{i}\}$ found upon the optimization and the corresponding $\Delta A$ constitute the training samples.
Network and optimization details.
The training was performed using a multi-layer perceptron (MLP) neural network for both approaches. The network takes as input a one-dimensional vector of attenuation difference and outputs molecular concentration changes.
We trained both networks with early stopping when they reached convergence. To find the optimal network architecture, we used the Ray Tune library [27] to validate different MLP architectures (width, number of hidden layers, and activation functions), learning rates, and batch sizes.
Least squares optimization for the Beer-Lambert law, excluding the scattering effect, can be performed via multiplication of the observed attenuation with the pseudoinverse of the absorption coefficients [5]. In order to perform the nonlinear least-squared optimization for the Beer-Lambert law model including scattering, we used the publicly available solver of the SciPy library [28]. We used the least-squares minimization obtained predictions as the ground truth to validate all the trained networks.
3 Results
3.1 Data.
For our experiments, we used two types of modified Beer-Lambert law formulation, with and without scattering, on two types of spectral datasets: broadband NIRS data for which the spectra were measured in light transmission mode [15], and hyperspectral data which were obtained in reflection mode [16].
Broadband NIRS.
The first dataset is composed of broadband NIRS spectra from a study analyzing 27 piglets’ brains in which a hypoxia state was induced [15]. The piglets were monitored for several hours, during which the carotid arteries were surgically isolated. This allowed controlling the blood flow to the brain. The details of the intervention protocol are described in [15]. The measurements contain around eight thousand spectra per piglet, of which we use the first thousand measurements (corresponding to the first $\sim$2.5 hours of measurement) where HI and recovery after HI are visible. As Eq. 2 requires defining a reference spectrum, analogously to [15], we used a spectrum at the very beginning of optical monitoring (i.e., before HI) for reference. We normalized the spectra with respect to dark noise and white reference. The normalized NIRS spectra before and after the intervention inducing hypoxia are shown in Fig. 3 (left), and predictions of the concentrations change over the course of the optical monitoring are shown in Fig. 3 (right). Out of the 27 piglets in the dataset, 25 had data available during HI, such that 19 were used for training, two for validation, and four for testing. For this dataset, we predict three types of molecules: oxyhemoglobin, deoxyhemoglobin, and differential cytochrome-c-oxidase (CCO) [^3] As the total CCO concentration may be assumed to not change within a few hours, the oxidized-reduced difference spectrum $\mu_{a}^{\text{diffCCO}}$ may be used to infer changes of both oxidized and reduced CCO [29]. , i.e. $\{\delta c_{i}\}=\{\delta c_{HbO_{2}},\delta c_{HHb},\delta c_{\text{diffCCO}}\}$, where $\delta c_{\text{diffCCO}}=\delta c_{oxCCO}-\delta c_{redCCO}$. We neglected the potential contribution to the spectra from water and fat due to their minimal change in concentrations during the monitoring and rather flat absorption spectra in the considered range of wavelengths between 740 and 900 nm [15, 30].
Figure 3: Optical spectra from the broadband NIRS study in [15] before and after inducing hypoxia in the piglet’s brain (left). On the right, predictions of the molecular concentration change $\{\delta c_{HbO_{2}},\delta c_{HHb},\delta c_{\text{diffCCO}}\}$ over the course of the optical monitoring (for the Beer-Lamber model without scattering).
Hyperspectral data.
The second dataset we used consists of hyperspectral data from the Helicoid project [16]. The Helicoid dataset comprises brain HSI images obtained in surgical conditions from 22 patients diagnosed with glioma. The HSI images provide a high spectral resolution of 826 bands spread between 400 and 1000 nm and a 2D spatial resolution of a few hundred pixels in each dimension. The images were also expert-annotated into three tissue classes: normal and tumor tissues, as well as blood vessels. Typical hyperspectral image and corresponding spectra are shown in Fig. 4. Different from the NIRS dataset, as a reference spectrum, we used a spectrum of the pixel belonging to the blood vessel class. We then subtracted the reference spectrum from all other spectra in the image. Besides predicting oxyhemoglobin $\delta c_{HbO_{2}}$ and deoxyhemoglobin $\delta c_{HHb}$, we again infer the differential cytochrome-c-oxidase concentration due to its role in capturing oxidative metabolic activity. We separately predicted oxidized cytochrome-c-oxidase and reduced cytochrome-c-oxidase, as the total CCO concentration may not be assumed to remain constant in space. We also predict water and fat since, for these molecules, one cannot assume minimal concentration change across different tissue types as in the case of the NIRS spectra. Fig. 5 showcases examples of molecular inference for the HSI images from the Helicoid dataset. Out of the nine patients with glioblastoma in the dataset, six with distinct class labeling were chosen, and three patients were used for training, one for validation, and two for testing. Note that patients might have multiple images taken, such that the training set consists of five, the validation set of one, and the test set of three images.
Figure 4: Optical spectra from the HSI study of patients diagnosed with glioma [16] for different tissue types: tumor, normal tissue, and blood vessels.
Figure 5: An example of HSI image from the Helicoid dataset [16] (left). Each pixel in the shown 2D image possesses a spectral signature with 826 bands. From this signature, we predict the molecular concentration change for hemodynamic $\delta c_{HbT}=\delta c_{HbO_{2}}+\delta c_{HHb}$ (middle), and metabolic $\delta c_{\text{diffCCO}}=\delta c_{oxCCO}-\delta c_{redCCO}$ characterization (right).
3.2 Experiments.
Scattering vs Non-scattering.
First, we test different Beer-Lambert formulations - with and without scattering. This test is motivated by our desire to assess whether networks can substitute not only linear solvers (without scattering) but also non-linear ones (with scattering) since the latter require a large computation time for optimization.
Figure 6: Comparison between predictions using linear (no scattering) and non-linear (with scattering) models for NIRS (left) and HSI (right) spectra. The inclusion of scattering into the formulation of the Beer-Lambert law notably improves the spectral fit to real data.
Fig. 6 shows spectral fits using both formulations. The model with scattering provides a clearly better fit. It allows us to better describe the peak around 760 nm for the NIRS data, while for the HSI data, the inclusion of scattering is often merely necessary for an accurate fit of the spectra. The necessity of the scattering consideration in the Beer-Lambert model for the HSI data can be explained by the more pronounced contribution of the scattering process  [^4] For the HSI data, we infer the difference in molecular composition between different spatial locations on an image, i.e., between different tissue types. The scattering property across brain tissues can significantly vary, and thus, the scattering shapes markedly the differential spectra. In contrast, for the NIRS data, we perform the differential spectroscopy analysis not in space but in time (comparing two spectra for the same location taken at different time points), meaning that the molecular inference is performed for the same tissue type..
Note, as a by-product of our analysis, we observe that the molecular inference driven by the physical model leads to better contrasting the blood vessel tree compared to the RGB image, Fig. 5 (middle), and capturing metabolic activity (right).
Synthetic vs Real.
i
ii
Figure 7: Comparison between inference of the molecular composition using the standard optimization methods and proposed network-based inference for training strategy (a) and (b) on both datasets. The top row compares strategy (a), left, with strategy (b), right, for the NIRS dataset. The bottom row compares the inference of the hemodynamic signal of the optimization-based result, left, with strategy (a), middle, and strategy (b), right. Despite the fact that the training distribution is different from the test distribution (especially for strategy (a) with synthetic data used for training), the network predictions are closely aligned with the predictions by the standard solver.
Fig. 7 demonstrates the results of the experiment in which we test the network trained on synthetic data collected according to strategy (a) and on real data according to strategy (b). The latter provides qualitatively closer fits for the shown piglet, which is also quantitatively supported as presented in table 1. Nevertheless, the predictions from both strategies are close to the ground truth predictions obtained by the least-squares optimizer, especially for NIRS data. Intuitively, this behavior is not within expectations as the spectra of synthetic data on which the network was trained and real data on which it was evaluated notably vary (e.g., due to noise). We attribute such behavior to the fact that the main spectral feature that networks learn to focus on is the global functional shape of the spectra. The local behavior of the spectral function is less informative, and thus, the local perturbations, i.e., random noise, do not affect the network performance significantly. This reasoning is further reinforced when comparing the network performance between NIRS and HSI data. For the HSI data, the noise has a pronounced wavelength dependency, in contrast to NIRS data (see Fig. 4), which results in the network performance degrading more notably compared to NIRS data.
Dataset Individual ID Spectral MAE Concentration MAE
Non-Scattering Scattering Strategy (a) Strategy (b)
507 0.012 2620.0122620.012\,2620.012 262 0.008 6010.0086010.008\,6010.008 601 0.013 5880.0135880.013\,5880.013 588 0.004 3180.0043180.004\,3180.004 318
Broadband 509 0.010 9150.0109150.010\,9150.010 915 0.009 3680.0093680.009\,3680.009 368 0.024 5700.0245700.024\,5700.024 570 0.004 8120.0048120.004\,8120.004 812
NIRS 511 0.008 0050.0080050.008\,0050.008 005 0.006 5950.0065950.006\,5950.006 595 0.011 5860.0115860.011\,5860.011 586 0.003 4080.0034080.003\,4080.003 408
512 0.011 9920.0119920.011\,9920.011 992 0.010 8820.0108820.010\,8820.010 882 0.012 5320.0125320.012\,5320.012 532 0.004 5910.0045910.004\,5910.004 591
012-01 0.032 6900.0326900.032\,6900.032 690 0.024 8960.0248960.024\,8960.024 896 0.173 4640.1734640.173\,4640.173 464 0.016 3950.0163950.016\,3950.016 395
HELICoiD 012-02 0.022 4110.0224110.022\,4110.022 411 0.021 9260.0219260.021\,9260.021 926 0.149 7520.1497520.149\,7520.149 752 0.025 8000.0258000.025\,8000.025 800
015-01 0.063 2640.0632640.063\,2640.063 264 0.025 3730.0253730.025\,3730.025 373 0.174 6690.1746690.174\,6690.174 669 0.015 3810.0153810.015\,3810.015 381
Table 1: Quantitative performance comparison of the different Beer-Lambert models and network training strategies on the test set of the two spectral datasets. To compare the two (non-scattering and scattering) models, we compute the mean absolute error of the spectral fit (denominated as ’Spectral MAE’) between the ground truth observed and predicted signals. The comparison of the two network training strategies is performed by assessing the mean absolute error of each strategy between the network and optimization-inferred concentrations (denominated as ’Concentration MAE’) of all considered chromophores. In the case of the HELICoiD dataset, only pixels labeled as normal, tumor, or blood were considered for these computations. The best-performing model and strategy for each individual is highlighted in bold.
Another explanation for this behavior can be that here, we try to learn with a neural network a solution to a linear system (or quasi-linear system, in the case of the model with scattering, where scattering contribution to the reflection spectra is minor compared to the absorption, as for NIRS data). We hypothesize that upon network training on the dataset representing solutions to linear systems, the network weights are learned to minimize the contribution from non-linear network units (e.g., activation functions), as such units are unnecessary to establish a linear mapping. Generally speaking, the non-linear units are a virtue and a curse of the networks. The virtue, because they allow to learn an arbitrary mapping, and the curse, because it is the very sequential application of non-linear units that blows up the predictions for input data unseen during training. But if the contribution from non-linear units of the network is minimized, we would not expect such blow-ups of the network’s predictions.
Within expectation were, however, the results obtained by the strategy (b). In contrast to training on the synthetic data, strategy (b) allows us to infer molecular composition almost identical to the convex optimization solver for both NIRS and HSI spectra.
Computational time.
Importantly, the proposed network-based optimization comes with a significant speed-up in computational time. In Fig. 8, we show a comparative analysis for performing chromophore composition inference using standard least-squares solvers (based on gradient update or pseudoinverse) and our proposed approach. As solving the linear system using the pseudoinverse requires the least amount of matrix multiplication, this method provides the fastest compute.
Figure 8: Comparison between inference time for various optimization approaches for varying number of spectra (from $10$ to $10^{5}$): including the pseudoinverse for the linear model (blue) and optimization-based (red) for the non-linear scattering model (both running on CPU), as well as network-based approach for scattering model running on CPU (orange) and GPU (green).
However, with the growth of the number of spectra for which we solve the optimization task, the matrix size for the inversion increases, and thus the computational time increases. Starting from ca. $10^{4}$ amount of spectra, the proposed network having a fixed amount of computational units becomes superior in terms of optimization time. Such runtime will remain approximately constant with a further increasing number of spectra, assuming sufficient GPU memory is available. More importantly, for non-linear systems, which are here represented as a Beer-Lambert model with the inclusion of scattering, one cannot utilize the pseudoinverse and has to resort to non-linear solvers like the ones based on gradient update. Such solvers are two to three orders of magnitude slower than the neural network approach, which has fixed compute time for linear and non-linear systems. As Fig. 8 shows, it takes ca. 0.7 ms for the network to infer biochemical composition for $10^{5}$ spectra on NVIDIA GeForce MX450 with 2048 MiB.
4 Conclusion
In the paper, we present proof of a data-driven concept for inferring molecular composition change from optical reflection spectra of brain tissue. We test the approach on various datasets (broadband NIRS and HSI data) and physical models (with and without scattering, i.e., linear and non-linear). We achieve a significant speed-up compared to traditional least-squares optimization, reaching subsecond time for inferring molecular composition. The subsecond timing stays the same for linear models (without scattering) and non-linear models (with scattering). Such timing would allow for real-time tissue characterization for both NIRS and HSI imaging.
Disclosures
The authors declare no financial conflict of interest.
Code, Data, and Materials Availability
In support of open science, all our code is available under https://github.com/HyperProbe/SpectraFit. The HSI data is publicly available by the HELICoiD project. The bNIRS data may be obtained upon request.
Acknowledgments
The HyperProbe consortium has received funding from the European Union’s Horizon Europe research under grant agreement No. 101071040. UCL is supported by UKRI grant No. 10048387.